# Entanglement, Trace Anomaly, and Confinement in QCD

### v3

@@ Abstract
We formulate confinement in QCD as an entropic surface phenomenon. Quark and gluon quantum information is localized on a transverse, entangling two‑sphere of radius $R_{\!EE}$; at this radius the QCD vacuum—partitioned by a hadron into interior and exterior regions—reaches its maximal entanglement entropy. Lattice‑QCD determinations of the scalar (trace) gravitational form factors fix both $R_{\!EE}$ and the transverse trace‑anomaly density $\rho_{h}(R_{\!EE})$, yielding a parameter‑free slope $c_{h}=8\pi^{2}R_{\!EE}^{2}\rho_{h}(R_{\!EE})$ and a mechanical entropy $S_{\!EE}(y)=c_{h}\,y$ that grows linearly with rapidity $y$. The entropy gradient $\partial_{R}S_{\!EE}$ changes sign at $R_{\!EE}$: it pushes colored degrees of freedom outward for $r<R_{\!EE}$ and pulls them inward for $r>R_{\!EE}$, thereby localizing them on the codimension‑2 entangling two‑sphere $\Sigma_{\perp}=S^{2}_{R_{\!EE}}$ (which, in the infinite‑momentum frame (IMF), projects onto the transverse plane)—the “information wall.” This provides a high‑energy (large‑$y$) entropic confinement diagnostic that complements—rather than replaces—Wilson’s area‑law criterion, which probes long‑distance dynamics near the rest frame ($y\!\to\!0$). Imposing unitarity on an entropic ansatz for the amplitude yields $\sigma(s)\!\propto\!y^{\delta}$. World data favor $\delta=2$ for elastic $pp\,(p\bar{p})$ scattering and heavy‑quark photoproduction, whereas $\phi$ photoproduction favors a softer $\delta=0.387$. All extracted cross sections remain well below the Froissart–Martin bound. These results provide a confinement criterion quantified directly from non‑perturbative QCD inputs, unifying the trace anomaly, entanglement entropy, and high‑energy scattering within a single quantitative framework.
Introduction.— Why do colored quarks and gluons never emerge from hadrons? The textbook answer invokes an area law for large Wilson loops, which diagnoses a linear confining potential in the IR [1]. Yet that criterion carries no information about ultraviolet (UV) dynamics, where asymptotic freedom reigns.
Here we propose a complementary, UV–sensitive diagnostic: entropic confinement. Quantum information carried by quarks and gluons localizes on a codimension‑2 entangling surface of fixed radius $R_{\!EE}$. Geometrically, the entangling surface is the two–sphere $\Sigma_{\perp}=S^{2}_{R}$ equipped with the induced metric $h_{ab}$ and area element $\sqrt{h}\,d^{2}x$. The universal contribution to the von Neumann entropy of the reduced state is then controlled by the QCD trace anomaly on $\Sigma_{\perp}$ [2, 3]:
@@
as established by general arguments in relativistic QFT [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. [see Eq. (S13) of Appendix for a detailed derivation of Eq. (1).] In the infinite‑momentum frame (IMF), the two‑sphere $S^{2}_{R}$ projects onto the transverse (impact‑parameter) plane, and the surface measure reduces to the standard light‑front impact‑parameter measure used to define parton densities [15, 16].We use this mapping solely to express densities; the underlying entangling surface remains $S^{2}_{R}$.
By “mechanical entropy” we mean the universal logarithmic term $S_{\!EE}^{\rm univ}$ evaluated under two physical renormalization conditions: (i) a maximal‑entropy radius $R=R_{\!EE}$ set by the hadron; and (ii) a Lorentz‑contracted cutoff $\epsilon=R_{\!EE}e^{-y}$ for boost rapidity $y$. With these choices, spherical symmetry on $\Sigma_{\perp}$ gives $\int_{\Sigma_{\perp}}\!\sqrt{h}\,\langle T^{\mu}{}_{\mu}\rangle=4\pi R_{\!EE}^{2}\,\rho_{h}(R_{\!EE})$, so that
@@
For complementary 1+1D derivations of the linear‑in‑$y$ growth of entanglement, see [17, 18, 19, 20]. In particular, Ref. [20] computes the energy-momentum tensor (EMT) of a “hadron‑like” primary of dimension $h$ and finds (their Eq. 3.15) in pure 1+1D CFT $T_{++}(x_{+})=C_{h}/x_{+}^{2}$ with $C_{h}=h$; in the mass‑deformed case (their Eq. 3.19) one reads off the same coefficient from $C_{h}=x_{+}^{2}T_{++}(x_{+})$ at $x_{+}\!\to\!0,\infty$. This gives a universal slope $S_{EE}^{\rm univ}(y)=2\pi\,C_{h}\,y$, implying the dictionary $C_{h}=c_{h}/(2\pi)$ when compared to Eq. (2). Our result is the 3+1D QCD realization: we determine $c_{h}$ non‑perturbatively from the hadron’s trace anomaly localized on the entangling two‑sphere, $c_{h}=8\pi^{2}R_{EE}^{2}\rho_{h}(R_{EE})$, using lattice‑QCD gravitational form factor (GFF) inputs in 2D impact-parameter space Eq. (9).
It is convenient to parameterize the slope $c_{h}$ in terms of an operational (non‑perturbative) occupation number on the information wall:
$\frac{c_{h}}{2\pi}\ \stackrel{{\scriptstyle\text{op.}}}{{=}}\ \int_{0}^{1}\!dx\int d^{2}k_{\perp}\int d^{2}b_{\perp}\,\delta_{\Sigma}(b_{\perp})\,n_{q+g}(x,\bm{k}_{\perp},\bm{b}_{\perp})\,$, (3)
where $\delta_{\Sigma}(b_{\perp})\equiv\frac{\delta(b_{\perp}-R_{\!EE})}{2\pi R_{\!EE}}$, and $n_{q+g}(x,\bm{k_{\perp}},\bm{b}_{\perp})$ denotes the total (sea‑quark + gluon) occupation number with longitudinal momentum fraction $x$ and transverse momentum $\bm{k_{\perp}}$ per unit transverse phase‑space area $d^{2}b_{\perp}$; this identification serves as a non‑perturbative analogue of the gluon occupation number in the CGC/BFKL framework [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. Also note that $c_{h}$ per nucleon shows rapid saturation with increasing mass number $A$ in nuclei; see Fig. 1.
Formally, we define the hadron’s entanglement by starting from the vacuum density operator $|0\rangle\!\langle 0|$ and tracing out all quark‑ and gluon‑field modes whose support lies outside a sphere of radius $R$. The reduced density matrix admits the Euclidean path‑integral representation [2]
@@
Percent-level lattice QCD calculations of the scalar (trace) gravitational form factors (GFFs) [39, 40, 41, 42] provide a direct, non-perturbative handle on $\langle T^{\mu}_{\ \mu}\rangle$. Independent experimental extractions corroborate these values: (i) the quark-sector GFF is obtained from deep-virtual Compton scattering at Jefferson Lab [43], analyzed with the dispersive–factorization approaches of [44, 45, 46]; (ii) the gluon-sector GFF is deduced from near-threshold $J/\psi$ photoproduction [47] using both holographic-QCD [48, 49] and an NRQCD–GPD factorization scheme [50].
The sign‑changing entropy gradient is an intrinsic property of the QCD vacuum, akin to Casimir-like stresses in non-Abelian gauge theories [51], and not a by‑product of the $2^{++}$$0^{++}$ resonances admixture that governs the Breit‑frame three‑dimensional EMT pressure inside a hadron [52, 53, 54, 55, 56, 57] (see Fig. S2 in the Appendix for the separate mechanical‑pressure node).
An entropic ansatz relates an elastic $2\!\to\!2$ cross section to the mechanical entropy of the color field produced in the collision [58, 59]; see Cross Sections and Data below for details. Ref. [60], following the exponential entropic ansatz of Ref. [19] for DIS structure functions, models total cross sections as $\sigma\propto e^{\delta y}$. Although this form reproduces vector–meson photoproduction data at high energies, it fails near threshold and overshoots the Froissart–Martin unitarity bound in the $y\!\to\!\infty$ limit. By contrast, our power‑law ansatz $\sigma\propto y^{\delta}$ describes both threshold and high‑energy photoproduction data with a single exponent and remains well below the Froissart–Martin limit across the full kinematic range, as illustrated in Figs. 4a4c. See also [61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73] for other proposals which relate entanglement entropy with multiplicity of final-state hadrons produced in high-energy collisions.
In the remainder of this Letter we (i) summarize the lattice‑QCD inputs that fix $\rho_{h}$ and $R_{\!EE}$; (ii) quantify the entropic force associated with $\partial_{R}S_{\!EE}$; and (iii) show how unitarity reduces an entropic ansatz for elastic $2\!\to\!2$ processes to $\sigma(s)=\mathcal{N}^{2}y^{\delta}$, with $\delta$ reflecting the UV sensitivity of the channel.
@@
Spin‑0 nuclei.— With $n_{A}=3A-1$,
@@
Summing over $i=q,g$ and evaluating at $b_{\perp}=R_{\!EE}$ fixes $c_{h}=8\pi^{2}R_{\!EE}^{2}\rho_{h}(R_{\!EE})$ in Eq. (2). All lattice fit parameters are given in Table S1 of Appendix.
@@
Figure 3: Visual comparison of proton radii using lattice gravitational form factor (5) from [40], see Table S1: maximum-entropy radii $R_{\!EE}$, mechanical core $R_{q,g}$ (S18) (from the node of the 3D Breit‑frame pressure density, see Fig. S2), rms scalar radii $r_{s}$ (S20), rms energy radii $r_{00}$ (S20), and the experimental rms charge radius $r_{c}$ from PRad [81]. All rms radii shown are computed separately for the quark and for the gluon sectors using 3D Breit-frame densities and Eq. (S20) in the Appendix; no q+gq\!+\!g sum is taken. However, $R_{\!EE}$ is computed from $q+g$ 2D (impact-parameter space) trace anomaly density Eq. (9).
@@
Mechanical Entanglement and Confinement.— Equation (2), together with Figs. 1 and 2, shows that the entropy gradient reverses sign at $r=R_{\!EE}$ (see Table 1 for values of $R_{\!EE}$ for pion and nuclei, also see Fig. 3 for a visual comparison of various proton radii inferred from the lattice GFFs). Accordingly, the radial derivative $\partial_{R}S_{\!EE}$ acts as an entropic pressure or force in the sense of Refs. [89, 90]. A parton that strays inside the surface ($r<R_{\!EE}$) is pushed outward, whereas one that moves outside ($r>R_{\!EE}$) is pulled back. Colored degrees of freedom are therefore locked onto the two-dimensional “information wall” at $r=R_{\!EE}$. This entropic bag complements the Wilson area‑law criterion: it provides a high‑energy (UV) diagnostic while the area law governs long‑distance (IR) dynamics.
@@
and, because the phase accumulates additively with rapidity, we take an entropy‑controlled eikonal $\chi(b_{\perp},y)\propto y$. This yields a single‑scale power law at high energy.
Consequently, for each elastic/exclusive channel, we write
@@
UV–dominated channels.— Rapidity is defined as $y=\operatorname{arcosh}\gamma=\ln\bigl(\gamma+\sqrt{\gamma^{2}-1}\bigr)$. For vector–meson photoproduction with $s=W^{2}=\left(p_{\gamma}+p_{p}\right)^{2}$, we adopt the dimensionless beam–energy ratio $\gamma_{\text{beam}}\equiv E_{\gamma}/E_{\gamma,\mathrm{thr}}(V)$, where $E_{\gamma,\mathrm{thr}}(V)=\bigl[(m_{N}+M_{V})^{2}-m_{N}^{2}\bigr]/(2m_{N})$, so that $\gamma_{\text{beam}}=1$ at threshold and $y_{\text{beam}}=\operatorname{arcosh}\gamma_{\text{beam}}$. For elastic $pp$ scattering (with $\sqrt{s}=W$) in the c.m. frame the Lorentz factor of either proton is $\gamma_{\text{CM}}\equiv\sqrt{s}/(2m_{N})$, yielding $y_{\text{CM}}=\operatorname{arcosh}\gamma_{\text{CM}}$. Reactions characterized by compact transverse sizes (elastic $pp$, $J/\psi$, and $\Upsilon$ production) sample the full ultraviolet content of the wall; their exponents saturate at the unitarity limit $\delta=2$:
@@
All extracted cross sections remain well below the canonical Froissart–Martin scale $(\pi/m_{\pi}^{2})\ln^{2}\!s$ for $\sigma_{\mathrm{tot}}$ (with $\pi/m_{\pi}^{2}\simeq 62.8~\mathrm{mb}$ after unit conversion): elastic $pp$ lies by factors $\gtrsim 2$ below at present energies, whereas heavy‑quarkonia photoproduction is suppressed by $\sim 10\!-\!10^{3}$.
Conclusions.— We have shown that the QCD trace anomaly, quantified with percent-level lattice determinations of the scalar gravitational form factors, fixes a transverse “information wall” of radius $R_{\!EE}$. Tracing over quark and gluon modes outside that surface yields a finite, parameter-free mechanical entropy $S_{\!EE}=c_{h}\,y$ with $c_{h}$ fixed entirely by lattice inputs—whose sign-changing radial derivative locks colored fields to the wall at all momentum scales. This entropic mechanism provides a high‑energy diagnostic that complements the traditional Wilson‑loop area law: it remains consistent with asymptotic freedom in the ultraviolet and with linear confinement in the infrared. Imposing unitarity on an entropic ansatz for the amplitude yields $\sigma(s)\propto y^{\delta}$. Hard processes ($pp$, $J/\psi$, $\Upsilon$) saturate the unitarity limit $\delta=2$, exhibiting the familiar $y^{2}\!\sim\!\ln^{2}\!s$ growth pattern, while the more diffuse $\phi$ channel selects $\delta_{\phi}=0.387$. All extracted cross sections remain well below the canonical Froissart–Martin bound $(\pi/m_{\pi}^{2})\ln^{2}\!s$ for $\sigma_{\mathrm{tot}}$.
Outlook.— The entropic framework developed here delivers a quantitative baseline against which future high-precision data can be benchmarked. Upcoming measurements at the HL-LHC, Jefferson Lab (CLAS12, GlueX, and SoLID), and the Electron–Ion Collider will confront its predictions across an unprecedented kinematic range. Any systematic deviation—for example from odderon exchange, Regge exchanges, exotic multiquark resonances—will stand out cleanly against the fully quantified entropy background, offering a sharpened path to discovering genuinely new QCD dynamics.

### v4

@@ Abstract
We formulate confinement in quantum chromodynamics (QCD) as an entropic surface phenomenon. Quark and gluon quantum information is localized on a transverse, entangling two-sphere of radius $R_{\!EE}$; at this radius the QCD vacuum—partitioned by a hadron into interior and exterior regions—reaches its maximal entanglement entropy. Lattice-QCD determinations of the scalar (trace) gravitational form factors fix both $R_{\!EE}$ and the transverse trace-anomaly density $\rho_{h}(R_{\!EE})$, yielding a parameter-free slope $c_{h}=8\pi^{2}R_{\!EE}^{2}\rho_{h}(R_{\!EE})$ and a mechanical entropy $S_{\!EE}(y)=c_{h}\,y$ that grows linearly with rapidity $y$. The entropy gradient $\partial_{R}S_{\!EE}$ changes sign at $R_{\!EE}$: it pushes colored degrees of freedom outward for $r<R_{\!EE}$ and pulls them inward for $r>R_{\!EE}$, thereby localizing them on the codimension-2 entangling two-sphere $\Sigma_{\perp}=S^{2}_{R_{\!EE}}$ (which, in the infinite-momentum frame (IMF), projects onto the transverse plane)—the “information wall.” This provides a high-energy (large-$y$) entropic confinement diagnostic that complements—rather than replaces—Wilson’s area-law criterion, which probes long-distance dynamics near the rest frame ($y\!\to\!0$). Imposing unitarity on an entropic ansatz for the amplitude yields $\sigma(s)\!\propto\!y^{\delta}$. World data favor $\delta=2$ for elastic $pp\,(p\bar{p})$ scattering and heavy-quark photoproduction, whereas $\phi$ photoproduction favors a softer $\delta=0.387$. All extracted cross sections remain well below the Froissart–Martin bound. These results provide a confinement criterion quantified directly from nonperturbative QCD inputs, unifying the trace anomaly, entanglement entropy, and high-energy scattering within a single quantitative framework.
Introduction.— Why do colored quarks and gluons never emerge from hadrons? The textbook answer invokes an area law for large Wilson loops, which diagnoses a linear confining potential in the infrared (IR) [1]. Yet that criterion carries no information about ultraviolet (UV) dynamics, where asymptotic freedom reigns.
Here we propose a complementary, UV–sensitive diagnostic: entropic confinement. Quantum information carried by quarks and gluons localizes on a codimension-2 entangling surface of fixed radius $R_{\!EE}$. Geometrically, the entangling surface is the two–sphere $\Sigma_{\perp}=S^{2}_{R}$ equipped with the induced metric $h_{ab}$ and area element $\sqrt{h}\,d^{2}x$. The universal contribution to the von Neumann entropy of the reduced state is then controlled by the quantum chromodynamics (QCD) trace anomaly on $\Sigma_{\perp}$ [2, 3]:
@@
as established by general arguments in relativistic QFT [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. See the Supplemental Material (included at the end of this arXiv version) for a detailed derivation of Eq. (1). In the infinite-momentum frame (IMF), the two-sphere $S^{2}_{R}$ projects onto the transverse (impact-parameter) plane, and the surface measure reduces to the standard light-front impact-parameter measure used to define parton densities [15, 16]. We use this mapping solely to express densities; the underlying entangling surface remains $S^{2}_{R}$.
By “mechanical entropy” we mean the universal logarithmic term $S_{\!EE}^{\rm univ}$ evaluated under two physical renormalization conditions: (i) a maximal-entropy radius $R=R_{\!EE}$ set by the hadron; and (ii) a Lorentz-contracted cutoff $\epsilon=R_{\!EE}e^{-y}$ for boost rapidity $y$. With these choices, spherical symmetry on $\Sigma_{\perp}$ gives $\int_{\Sigma_{\perp}}\!\sqrt{h}\,\langle T^{\mu}{}_{\mu}\rangle=4\pi R_{\!EE}^{2}\,\rho_{h}(R_{\!EE})$, so that
@@
For complementary 1+1D derivations of the linear-in-$y$ growth of entanglement, see [17, 18, 19, 20]. In particular, Ref. [20] computes the energy-momentum tensor (EMT) of a “hadron-like” primary of dimension $h$ and finds (their Eq. 3.15) in pure 1+1D CFT $T_{++}(x_{+})=C_{h}/x_{+}^{2}$ with $C_{h}=h$; in the mass-deformed case (their Eq. 3.19) one reads off the same coefficient from $C_{h}=x_{+}^{2}T_{++}(x_{+})$ at $x_{+}\!\to\!0,\infty$. This gives a universal slope $S_{EE}^{\rm univ}(y)=2\pi\,C_{h}\,y$, implying the dictionary $C_{h}=c_{h}/(2\pi)$ when compared to Eq. (2). Our result is the 3+1D QCD realization: we determine $c_{h}$ nonperturbatively from the hadron’s trace anomaly localized on the entangling two-sphere, $c_{h}=8\pi^{2}R_{EE}^{2}\rho_{h}(R_{EE})$, using lattice-QCD gravitational form factor (GFF) inputs in 2D impact-parameter space Eq. (9).
It is convenient to parametrize the slope $c_{h}$ in terms of an operational (nonperturbative) occupation number on the information wall:
$\frac{c_{h}}{8\pi^{2}R_{\!EE}^{2}\,}\ \stackrel{{\scriptstyle\text{op.}}}{{=}}\ \int_{0}^{1}\!dx\int d^{2}k_{\perp}\int d^{2}b_{\perp}\,\delta_{\Sigma}(b_{\perp})\,n_{q+g}(x,\bm{k}_{\perp},\bm{b}_{\perp})\,$, (3)
where $\delta_{\Sigma}(b_{\perp})\equiv\frac{\delta(b_{\perp}-R_{\!EE})}{2\pi R_{\!EE}}$, and $n_{q+g}(x,\bm{k_{\perp}},\bm{b}_{\perp})$ denotes the total (sea-quark + gluon) occupation number with longitudinal momentum fraction $x$ and transverse momentum $\bm{k_{\perp}}$ per unit transverse phase-space area $d^{2}b_{\perp}$; this identification serves as a nonperturbative analog of the gluon occupation number in the CGC/BFKL framework [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. Also note that $c_{h}$ per nucleon shows rapid saturation with increasing mass number $A$ in nuclei; see Fig. 1.
Formally, we define the hadron’s entanglement by starting from the vacuum density operator $|0\rangle\!\langle 0|$ and tracing out all quark- and gluon-field modes whose support lies outside a sphere of radius $R$, see Fig. S1 in the Supplemental Material. The reduced density matrix admits the Euclidean path-integral representation [2]
@@
Percent-level lattice QCD calculations of the scalar (trace) gravitational form factors (GFFs) [39, 40, 41, 42] provide a direct, nonperturbative handle on $\langle T^{\mu}_{\ \mu}\rangle$. Independent experimental extractions corroborate these values: (i) the quark-sector GFF is obtained from deep-virtual Compton scattering at Jefferson Lab [43], analyzed with the dispersive–factorization approaches of [44, 45, 46]; (ii) the gluon-sector GFF is deduced from near-threshold $J/\psi$ photoproduction [47] using both holographic-QCD [48, 49] and an NRQCD–GPD factorization scheme [50].
The sign-changing entropy gradient (see Fig. 2) is an intrinsic property of the QCD vacuum, akin to Casimir-like stresses in non-Abelian gauge theories [51], and not a by-product of the $2^{++}$$0^{++}$ resonances admixture that governs the Breit-frame three-dimensional EMT pressure inside a hadron [52, 53, 54, 55, 56, 57] (see Fig. S2 in the Supplemental Material for the separate mechanical-pressure node).
An entropic ansatz relates an elastic $2\!\to\!2$ cross section to the mechanical entropy of the color field produced in the collision [58, 59]; see Cross sections and data below for details. Ref. [60], following the exponential entropic ansatz of Ref. [19] for DIS structure functions, models total cross sections as $\sigma\propto e^{\delta y}$. Although this form reproduces vector-meson photoproduction data at high energies, it fails near threshold and overshoots the Froissart–Martin unitarity bound in the $y\!\to\!\infty$ limit. By contrast, our power-law ansatz $\sigma\propto y^{\delta}$ describes both threshold and high-energy photoproduction data with a single exponent and remains well below the Froissart–Martin limit across the full kinematic range, as illustrated in Figs. 4a4c. See also [61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73] for other proposals which relate entanglement entropy with multiplicity of final-state hadrons produced in high-energy collisions.
In the remainder of this paper we (i) summarize the lattice-QCD inputs that fix $\rho_{h}$ and $R_{\!EE}$; (ii) quantify the entropic force associated with $\partial_{R}S_{\!EE}$; and (iii) show how unitarity reduces an entropic ansatz for elastic $2\!\to\!2$ processes to $\sigma(s)=\mathcal{N}^{2}y^{\delta}$, with $\delta$ reflecting the UV sensitivity of the channel.
@@
Spin-0 nuclei.— With $n_{A}=3A-1$,
@@
Summing over $i=q,g$ and evaluating at $b_{\perp}=R_{\!EE}$ fixes $c_{h}=8\pi^{2}R_{\!EE}^{2}\rho_{h}(R_{\!EE})$ in Eq. (2). All lattice fit parameters are summarized in Table S1 of the Supplemental Material.
@@
Figure 3: Visual comparison of proton radii using lattice gravitational form factor (5) from [40], see Table S1: maximum-entropy radius $R_{\!EE}$, mechanical core $R_{q,g}$ (S18) (from the node of the 3D Breit-frame pressure density, see Fig. S2), rms scalar radii $r_{s}$ (S20), rms energy radii $r_{00}$ (S20), and the experimental rms charge radius $r_{c}$ from PRad [81]. All rms radii shown are computed separately for the quark and for the gluon sectors using 3D Breit-frame densities and Eq. (S20) in the Supplemental Material; no q+gq\!+\!g sum is taken. However, $R_{\!EE}$ is computed from $q+g$ 2D (impact-parameter space) trace anomaly density Eq. (9).
@@
Mechanical entanglement and confinement.— Equation (2), together with Figs. 1 and 2, shows that the entropy gradient reverses sign at $r=R_{\!EE}$ (see Table 1 for values of $R_{\!EE}$ for pion and nuclei, also see Fig. 3 for a visual comparison of various proton radii inferred from the lattice GFFs). Accordingly, the radial derivative $\partial_{R}S_{\!EE}$ acts as an entropic pressure or force in the sense of Refs. [89, 90]. A parton that strays inside the surface ($r<R_{\!EE}$) is pushed outward, whereas one that moves outside ($r>R_{\!EE}$) is pulled back. Colored degrees of freedom are therefore locked onto the two-dimensional “information wall” at $r=R_{\!EE}$. This entropic bag complements the Wilson area-law criterion: it provides a high-energy (UV) diagnostic while the area law governs long-distance (IR) dynamics.
@@
and, because the phase accumulates additively with rapidity, we take an entropy-controlled eikonal $\chi(b_{\perp},y)\propto y$. This yields a single-scale power law at high energy. In particular, for small-eikonal (“moderate opacity”) regime, we find $\sigma_{\mathrm{el}}(s)\propto\int d^{2}b\,|\chi|^{2}\propto y^{2}$.
Consequently, for each elastic/exclusive channel, we write the entropic ansatz
@@
UV–dominated channels.— Rapidity is defined as $y=\operatorname{arcosh}\gamma=\ln\bigl(\gamma+\sqrt{\gamma^{2}-1}\bigr)$. For vector-meson photoproduction with $s=W^{2}=\left(p_{\gamma}+p_{p}\right)^{2}$, we adopt the dimensionless beam–energy ratio $\gamma_{\text{beam}}\equiv E_{\gamma}/E_{\gamma,\mathrm{thr}}(V)$, where $E_{\gamma,\mathrm{thr}}(V)=\bigl[(m_{N}+M_{V})^{2}-m_{N}^{2}\bigr]/(2m_{N})$, so that $\gamma_{\text{beam}}=1$ at threshold and $y_{\text{beam}}=\operatorname{arcosh}\gamma_{\text{beam}}$. For elastic $pp$ scattering (with $\sqrt{s}=W$) in the c.m. frame the Lorentz factor of either proton is $\gamma_{\text{CM}}\equiv\sqrt{s}/(2m_{N})$, yielding $y_{\text{CM}}=\operatorname{arcosh}\gamma_{\text{CM}}$. Reactions characterized by compact transverse sizes (elastic $pp$, $J/\psi$, and $\Upsilon$ production) sample the full ultraviolet content of the wall; their exponents saturate at the unitarity limit $\delta=2$:
@@
All extracted cross sections remain well below the canonical Froissart–Martin scale $(\pi/m_{\pi}^{2})\ln^{2}\!s$ for $\sigma_{\mathrm{tot}}$ (with $\pi/m_{\pi}^{2}\simeq 62.8~\mathrm{mb}$ after unit conversion): elastic $pp$ lies by factors $\gtrsim 2$ below at present energies, whereas heavy-quarkonia photoproduction is suppressed by $\sim 10\!-\!10^{3}$.
Conclusions.— We have shown that the QCD trace anomaly, quantified with percent-level lattice determinations of the scalar gravitational form factors, fixes a transverse “information wall” of radius $R_{\!EE}$. Tracing over quark and gluon modes outside that surface yields a finite, parameter-free mechanical entropy $S_{\!EE}=c_{h}\,y$ with $c_{h}$ fixed entirely by lattice inputs—whose sign-changing radial derivative locks colored fields to the wall at all momentum scales. This entropic mechanism provides a high-energy diagnostic that complements the traditional Wilson-loop area law: it remains consistent with asymptotic freedom in the ultraviolet and with linear confinement in the infrared. Imposing unitarity on an entropic ansatz for the amplitude yields $\sigma(s)\propto y^{\delta}$. Hard processes ($pp$, $J/\psi$, $\Upsilon$) saturate the unitarity limit $\delta=2$, exhibiting the familiar $y^{2}\!\sim\!\ln^{2}\!s$ growth pattern, while the more diffuse $\phi$ channel selects $\delta_{\phi}=0.387$. All extracted cross sections remain well below the canonical Froissart–Martin bound $(\pi/m_{\pi}^{2})\ln^{2}\!s$ for $\sigma_{\mathrm{tot}}$.
Outlook.— The entropic framework developed here delivers a quantitative baseline against which future high-precision data can be benchmarked. Upcoming measurements at the HL-LHC, Jefferson Lab (CLAS12, GlueX, and SoLID), and the Electron-Ion Collider will confront its predictions across an unprecedented kinematic range. Any systematic deviation—for example from odderon exchange, Regge exchanges, exotic multiquark resonances—will stand out cleanly against the fully quantified entropy background, offering a sharpened path to discovering genuinely new QCD dynamics.